# Under-Canopy Navigation using Aerial Lidar Maps

### v1

@@ Abstract
Navigation, Field Robots, Robotics and Automation in Agriculture and Forestry
@@ I Introduction
Figure 1: Sensors mounted on a ground robot can suffer from limited range and occlusions. This local view often precludes efficient long-distance navigation in complex scenarios, leading to dead-ends or unsafe paths (bottom figure: yellow arrows). Aerial lidar data (top left figure) acquired prior to operation can provide meaningful information for global planning (top right and bottom figures: magenta paths) allowing efficient navigation.
@@
We address these problems in a novel framework for under-canopy navigation in forests assisted by above-canopy aerial lidar scans. Our system builds a 3D probabilistic occupancy map from overhead lidar observations and uses occupancy probabilities to estimate the likelihood of ground-level obstruction throughout the map. Standard occupancy mapping techniques approach the problem as a pure mapping task with perfect sensor poses. Ignoring the inevitable error in aerial sensor’s pose estimates (produced by, e.g., pose-graph optimisation) can result in poor mapping accuracy, especially at longer ranges in aerial scenarios (lidar hit-point position error increases approximately linearly with the product of angle and range). To address this issue, our mapping system explicitly takes into account pose-estimation uncertainty and averages occupancy probabilities over the set of possible aerial vehicle trajectories using Monte Carlo sampling. The aerial map is then used to plan a safe and efficient path for the ground robot before its deployment. We introduce two path planning cost functions designed to strike a balance between minimising the travel distance and avoiding potential obstructions (inferred from the occupancy map). Our path planner is based on D* Lite [3] which supports efficient replanning in case the ground robot encounters a previously unseen obstacle along its current path. Finally, we evaluate the performance of our full system and its components in extensive experiments conducted in simulated and real complex forest environments. In summary, the main contributions of this work are the following:
  1. 1.
    A framework for navigation in dense forest environments using aerial lidar observations as prior information;
  2. 2.
    A mapping algorithm to estimate the obstructed areas under the canopy using airborne lidar data;
  3. 3.
    A Monte Carlo sampling method to propagate uncertainty from the aerial sensor trajectory to 3D occupancy maps;
  4. 4.
    Experimental evaluation in simulation and real-world experiments in a challenging forest area using a mobile tracked robot.
@@ II Related Work
II-A Mapping using aerial sensing modalities
Navigation in unstructured environments has drawn attention recently [4, 5, 6], however, the utilisation of aerial sensing data to support ground navigation is a less explored idea. To create traversability maps used in global planning, Vandapel et al. [7] first segment ground points from airborne lidar, then calculate costs based on the “vegetationess” metric (i.e., ratio of ground points and total number of points falling in a 2D grid region), and a vehicle mobility analysis above the terrain. Alternatively, a combination of geometric features (e.g., ground elevations) and semantic terrain classes from airborne sensing data (lidar, color and multispectral image) is used to estimate traversability costs [8]. In a similar fashion, [9] combines semantic terrain classes of earth observation data, geometric features (e.g., slope) and also proprioceptive locomotion data (e.g., velocities, IMU) of different robotic platforms to generate traversal costs conditioned on the platform type and terrain class. Lastly, Stentz et al. [10] learn high-cost features in a self-supervised fashion by fusing air-ground sensing data (satellite image and ground-view lidar).
Despite their relevance, [7, 8, 10] neglect free-space information and [7, 8] measurement uncertainty in their costmap definitions, which are crucial elements in determining traversable areas. Our approach relies on explicit probabilistic reasoning over voxel occupancy (free, occupied, or unknown), while also accounting for sensor noise and pose-estimation error. While we focus our method on obstructions inferred from occupancy maps, complementary notions of traversability from learning methods (e.g., semantic-based costs) could also be fused in our framework.
II-B Sensor pose uncertainty in occupancy maps
@@
Standard occupancy mapping approaches [11] and many of its variants [12, 13, 14] all assume that sensor poses are perfectly known at the time lidar measurements are acquired. However, this approximation can easily lead to high occupancy errors if localisation uncertainty increases. To address this problem, O’Callaghan et al. [15] propose to incorporate sensor and localisation uncertainty in Gaussian Process (GP) occupancy mapping by considering inputs as noisy observations drawn from a Gaussian distribution, and integrating them in the kernel function. However, such modifications increase the computational cost of the approach by an order of magnitude, which may limit its application in large datasets even in offline settings.
@@
In this work, we revisit and extend the idea of using Monte Carlo (MC) sampling to incorporate localisation uncertainty in 2D occupancy maps. This approach is first proposed in [16] and is used more recently in [17]. The former integrates out poses from occupancy distributions using sample trajectories obtained from a particle filter, whereas the later performs the MC integration in the log-odds occupancy ratio. While we also marginalise out poses from occupancy probabilities using MC sampling, unlike [16, 17] our approach generates samples from the approximated posterior of a modern pose-graph SLAM system. Furthermore, we apply this idea to 3D mapping and conduct extensive experimental evaluations in simulated forest environments.
II-C Path planning
@@
Graph-based (D*, A*) or sampling-based (Rapidly Exploring Random Trees - RRT*) methods have been broadly used for global planning in costmaps [18]. When applied to occupancy maps, occupancy probabilities are often thresholded to classify grid cells as obstacle or free space, subsequently assigning high and low costs to them [19, 20]. Other works [7, 21] propose heuristic functions of terrain features to build costmaps used in the referred planners. However, planning on thresholded occupancy maps or feature-based costs without uncertainty information can create a false sense of confidence in the generated path, ultimately leading to long detours.
Banfi et al. [22] address the problem of planning under uncertainty by generating multiple path hypotheses in unknown areas and reasoning on locally acquired information from a “next-best view” (NBV) pose to take the best route.
@@
Unlike the above approaches, our system generates a minimum-cost path to the goal by minimising cost functions that blend obstruction risk with path length. Additionally, we use D* Lite to efficiently replan a new optimal path each time the ground robot faces a previously unmapped obstacle along its current path.
@@ III System Overview
Fig. 2 summarises our proposed system. The aerial mapping (Sec. IV) starts aggregating airborne lidar scans and estimating the sensor trajectory using a localisation system, namely Wildcat SLAM [24]. Then, we estimate the ground height from the point cloud (Sec. IV-A), in addition to building a 3D probabilistic occupancy map from the aerial lidar observations (Sec. IV-B). For improved map accuracy, we also encode the uncertainty from sensor poses into occupancy distributions by means of a Monte Carlo integration method. Next, the ground height estimates and occupancy values are fed into our scoring function to estimate the obstructed (i.e., non-traversable) areas over the terrain (Sec. IV-C). Obstruction scores are finally integrated in two evaluated cost functions, expected cost and log-reachability cost, ultimately used by D* Lite to generate optimal paths.
@@ IV-A Ground Filtering
The ground filtering module estimates the ground height from 3D point clouds, producing a 2D grid map representation of the terrain’s supporting surface, where each grid square cell (side: $0.25$ m) stores the height information. Height estimates are further used in our mapping approach as references during the estimation of obstructed regions above the terrain.
We first apply the Cloth Simulation Filtering (CSF) [25] to segment ground points from the global point cloud. CSF vertically ($Z$) inverts the point cloud data and simulates gravity acting on a semi-rigid cloth – modeled as interconnected mass-spring nodes – covering the inverted surface. Then, CSF segments ground points by retaining the closest points in the cloud to the CSF nodes within a height threshold in each grid cell. For cells without ground points, we use the closest CSF node as a virtual ground point. Finally, the minimum height point in each grid cell is used as the ground height estimate. By operating on inverted point clouds, CSF is less susceptible to segmentation errors (e.g., lidar canopy reflections or high vegetation mistakenly classified as ground) and generalises well to distinct environments using only few parameters.
@@ IV-B Occupancy mapping with imperfect pose estimates
In addition to the ground elevations, occupancy estimates of the environment are also used as inputs to the obstruction estimation module. We process lidar points above ground using the Occupancy Homogeneous Mapping (OHM) [26] to build an occupancy grid map. OHM estimates an occupancy probability for each 3D grid element (voxel) by taking into account lidar hits and pass-throughs in each voxel. Standard occupancy mapping methods such as OHM assume lidar poses are perfectly known. However, in practice, poses are estimated by other state estimation systems (e.g., pose-graph optimisation, or some other form of SLAM) which are inevitably subject to estimation error due to sensor noise. Ignoring pose estimation errors leads to highly inaccurate maps. We address this challenge by incorporating the estimation uncertainty of the aerial vehicle’s trajectory provided by the pose-estimation system into occupancy mapping.
The occupancy mapping module estimates the occupancy probability of each voxel given aerial lidar measurements. We denote the map by $\{p(m_{i}\mid\mathcal{Z})\}_{i}$ where $m_{i}$ is a Bernoulli random variable indicating the occupancy of the $i$th voxel, and $\mathcal{Z}$ denotes the set of all measurements from the aerial vehicle (e.g. lidar, IMU, and GNSS). This probability can be computed as follows:
@@
where $\mathcal{X}$ denotes the robot’s trajectory (represented by a finite set of poses). The first integrand $p(m_{i}\,|\,\mathcal{X},\mathcal{Z})$ in (1) is the voxel occupancy probability computed by standard occupancy mapping systems given trajectory $\mathcal{X}$. Note that voxel occupancy $m_{i}$ is conditionally independent from IMU and GNSS measurements given the trajectory $\mathcal{X}$. Thus one can further simplify this term as $p(m_{i}\,|\,\mathcal{X},\mathcal{Z}_{\text{lidar}})$ where $\mathcal{Z}_{\text{lidar}}\subset\mathcal{Z}$ denotes the set of lidar measurements. The second integrand $p(\mathcal{X}\,|\,\mathcal{Z})$ is the posterior distribution over robot trajectory given all measurements. This term is computed by the underlying SLAM system (e.g., Wildcat’s pose-graph optimisation module [24]). We use Monte Carlo integration to approximate the intractable integral in (1). This is done by sampling $N$ trajectories $\{\mathcal{X}^{[k]}\}_{k=1}^{N}$ from the pose-graph SLAM posterior $p(\mathcal{X}\mid\mathcal{Z})$ and computing the following expression,
$p(m_{i}\mid\mathcal{Z})\approx\frac{1}{N}\sum_{k=1}^{N}p(m_{i}\mid\mathcal{X}^{[k]},\mathcal{Z}).$ (2)
Equation (2) suggests the following scheme for incorporating pose-estimation uncertainty into occupancy mapping: (i) build an occupancy map using any standard occupancy mapping method such as OHM for each trajectory sample $\mathcal{X}^{[k]}$; and (ii) for each voxel compute the average of the resulting $N$ predicted occupancy probabilities. We sample trajectories from the marginal covariances of individual pose posteriors. After solving the pose-graph SLAM optimisation, such covariances can be approximated by leveraging the relevant diagonal blocks of the information matrix representing the full posterior $p(\mathcal{X}\mid\mathcal{Z})$. Hereafter we use the term expected occupancy map to refer to the result of this module.
@@ IV-C Obstruction score estimation
Figure 3: We reason on free, occupied and unknown voxels by integrating lidar measurements in an occupancy grid map. Later the occupancy probabilities are fused in our scoring method to estimate the final obstruction map.
We use our 3D expected occupancy map and the estimated ground height to assign a simple obstruction risk score to each ground-level 2D cell. This score is computed for each vertical column of $n$ 3D voxels starting from the ground cell as illustrated in Fig. 3. Formally, the score function $b:\mathbb{Z}\times\mathbb{Z}\to[0,1]$ assigns a value between zero and one to each 2D cell (i.e., the grid $XY$ coordinates of the corresponding column in the world frame). A larger value of $b(s)$ for a cell $s\triangleq(x,y)$ indicates a higher risk of obstruction at that cell. The number $n$ of column voxels is chosen according to the clearance height ($h$ in Fig. 3) required by the specified robot to pass through a column. Therefore, taller robots require using a larger $n$.
A cell $s$ might be considered as obstructed if at least one of the voxels in the corresponding column is occupied. This choice leads to $b(s)=1-\prod_{i\in\mathcal{I}_{s}}p(m_{i}=0\mid\mathcal{Z})$. However, because of the complexity of forest environments and the sparsity of aerial lidar rays reaching the ground level, the resulting score function will likely fail to correctly capture traversable cells where only the first voxel above ground is occupied by grass or an overhanging foliage is occupying the top voxel in the stack. An effective heuristic estimate is to use instead the weighted average of voxel occupancy probabilities within a column as the corresponding obstruction score, where the (fixed) weights reflect the significance of voxel occupancy. The resulting score function $b$ is given by
@@
where $\mathcal{I}_{s}$ denotes the set of indices of the $n$ voxels that are part of the vertical stack at cell $s$, and $p(m_{i}=1\mid\mathcal{Z})$ is the (expected) occupancy probability of voxel $i$ given the aerial measurements. Specifically, in our work $n=4$ which allows for a robot height of 111\,1m using a 0.250.250.25\,0.25m voxel size, and the voxels weights $w_{i}$ from bottom to top are set to $(1,2,2,2)$. The rationale behind these values is to assign a smaller weight to the first voxel above ground level since the tracked mobile robot used in our experiments can better navigate obstacles immediately above the ground (e.g., small rocks, grass).

### v2

@@ Abstract
Field Robots; Robotics and Automation in Agriculture and Forestry; Mapping
© 2024 IEEE. The paper will appear in the IEEE Robotics and Automation Letters (RA-L). Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
We address these problems in a novel framework for under-canopy navigation in forests assisted by above-canopy aerial lidar scans. Our system builds a 3D probabilistic occupancy map from overhead lidar observations and uses occupancy probabilities to estimate the likelihood of ground-level obstruction throughout the map. Standard occupancy mapping techniques approach the problem as a pure mapping task with perfect sensor poses. However, ignoring the inevitable error in aerial sensor’s pose estimates (produced by, e.g., pose-graph optimisation) can result in poor mapping accuracy, especially at longer ranges in aerial scenarios (lidar hit-point position error increases approximately linearly with the product of angle and range). To address this issue, our mapping system explicitly takes into account pose-estimation uncertainty and averages occupancy probabilities over the set of possible aerial vehicle trajectories using Monte Carlo sampling. The aerial map is then used to plan a safe and efficient path for the ground robot before its deployment. We introduce two path planning cost functions designed to strike a balance between minimising the travel distance and avoiding potential obstructions (inferred from the occupancy map). Our path planner is based on D* Lite [3] which supports efficient replanning in case the ground robot encounters a previously unseen obstacle along its current path. Finally, we evaluate the performance of our full system and its components in extensive experiments conducted in simulated and real complex forest environments.
Figure 1: Sensors mounted on a ground robot can suffer from limited range and occlusions. This local view often precludes efficient long-distance navigation in complex scenarios, leading to dead-ends or unsafe paths (right figure: yellow arrows). Aerial lidar data (top left figure) acquired prior to operation can provide meaningful information for global planning (bottom left and right figures: magenta paths) allowing efficient navigation.
In summary, the main contributions of this work are: 1) A method for navigation in dense forest environments using aerial lidar observations as prior information; 2) A mapping algorithm to estimate the obstructed areas under the canopy using airborne lidar data; 3) A Monte Carlo sampling method to propagate uncertainty from the aerial sensor trajectory to 3D occupancy maps; 4) Experimental evaluation in simulation and real-world experiments in a challenging forest area using a mobile tracked robot.
@@ II Related Work
A. Aerial mapping for ground navigation - Prior work has proposed a range of approaches for generating traversability or cost maps for ground navigation from aerial data. Applied to forested areas, Vandapel et al. [4] estimate costs from airborne lidar based on the “vegetationess” metric (ratio of ground points and total number of points falling in a 2D grid region) and a vehicle mobility analysis. Combinations of geometric features (e.g., ground elevation) and semantic terrain classes from airborne data (lidar, colour and multispectral images) are used to estimate traversability or binary costs in [5, 1]. Similarly, in [6] semantic terrain classes, geometric features (e.g., slope) and proprioceptive locomotion data (e.g., velocities, IMU) of different robotic platforms are processed to generate traversal costs conditioned on the platform type and terrain class. Lastly, Stentz et al. [7] learn high-cost features from satellite images using ground-view lidar features for self-supervision. Despite their relevance, these works neglect free-space information[4, 5, 7] and measurement uncertainty[1, 4, 5] in their costmap definitions, which are crucial elements for traversability prediction. Our approach relies on explicit probabilistic reasoning over voxel occupancy (free, occupied, or uncertain), while also accounting for sensor noise and pose-estimation error. We also incorporate probabilistic obstruction and path length in our proposed cost functions for planning with uncertain maps, unlike previous methods.
@@
B. Sensor pose uncertainty in occupancy maps - Standard occupancy mapping approaches [8] and many of its variants [9, 10, 11] all assume that sensor poses are perfectly known at the time lidar measurements are acquired. However, this approximation can easily lead to high occupancy errors if localisation uncertainty increases. To address this problem, it is possible to incorporate sensor and localisation uncertainty in Gaussian process (GP) occupancy mapping by considering inputs as noisy observations drawn from a Gaussian distribution and integrating them in the kernel function [12]. However, such modifications increase the computational cost of the approach by an order of magnitude, which may limit its application in large datasets even in offline settings.
@@
We revisit and extend the idea of using Monte Carlo (MC) sampling to incorporate localisation uncertainty in 2D occupancy maps. This approach was first proposed in [13] and used more recently in [14]. The former samples trajectories from a particle filter to marginalise poses from occupancy distributions, whereas the latter performs MC integration in the log-odds occupancy ratio. While we also marginalise out poses from occupancy probabilities using MC sampling, unlike [13, 14] our approach generates samples from the approximated posterior of a modern pose-graph SLAM system. Furthermore, we apply this idea to 3D mapping and conduct extensive experimental evaluations in simulated forest environments.
@@
C. Path planning - Graph-based (D*, A*) or sampling-based (Rapidly Exploring Random Trees - RRT*) methods have been broadly used for global planning in costmaps [15]. When applied to occupancy maps, occupancy probabilities are often thresholded to classify grid cells as obstacle or free space, subsequently assigning high and low costs to them [16, 17]. Other works [4, 18] propose heuristic functions of terrain features to build costmaps. However, planning on thresholded occupancy maps or feature-based costs without uncertainty information can create a false sense of confidence in the generated path, potentially leading to long detours. Some authors address planning under uncertainty by minimising a risk metric over map predictions [19]. In contrast, Banfi et. al [20] generate multiple path hypotheses in unknown areas, further reasoning on locally acquired information from a “next-best view” (NBV) pose to take the best route.
@@
Unlike the above approaches, our system generates a minimum-cost path to the goal by minimising cost functions that blend obstruction risk with path length. Additionally, we use D* Lite to efficiently and optimally replan when the ground robot encounters previously unmapped obstacles along its current path.
@@ III System Overview
Fig. 2 summarises our system. The aerial mapping (Sec. IV) aggregates airborne lidar scans and estimates the sensor trajectory using a localisation system (Wildcat SLAM [22]). Then, we estimate the ground height from the point cloud (Sec. IV-A), in addition to building a 3D probabilistic occupancy map from the aerial lidar observations (Sec. IV-B). For improved map accuracy, we also encode the uncertainty from sensor poses into occupancy distributions employing a Monte Carlo integration method. Next, the ground height estimates and occupancy values are fed into our scoring function to estimate the obstructed (non-traversable) areas over the terrain (Sec. IV-C). Obstruction scores are finally integrated in two evaluated cost functions, expected cost and log-reachability cost, ultimately used by D* Lite to generate optimal paths.
@@ IV-A Ground Filtering
The ground filtering module takes in a 3D point cloud and produces a gridded ($0.25$ m side-length squares) 2D height map representing the terrain’s supporting surface. Height estimates are used in our mapping approach to estimate obstructed regions above the terrain.
We first apply the Cloth Simulation Filtering (CSF) [23] to segment ground points from the global point cloud. CSF vertically ($Z$) inverts the point cloud data and simulates gravity acting on a semi-rigid cloth – modeled as interconnected mass-spring nodes – covering the inverted surface. Then, CSF segments ground points by retaining the closest points in the cloud to the CSF nodes within a height threshold ($h_{cc}$) in each grid cell. For cells without ground points, we use the closest CSF node as a virtual ground. Finally, the minimum height point in each grid cell is used as the ground height estimate. By operating on inverted point clouds, CSF is less susceptible to segmentation errors (e.g., lidar canopy reflections or high vegetation mistakenly classified as ground) and generalises well to distinct environments using only few parameters, i.e., cloth rigidness ($RI=3$) and height threshold ($h_{cc}=0.20m$).
@@ IV-B Occupancy mapping with imperfect pose estimates
In addition to the ground elevations, occupancy estimates of the environment are also input to the obstruction estimation module. We process lidar points above ground using Occupancy Homogeneous Mapping (OHM) [24] to build an occupancy grid map. OHM estimates occupancy probability by taking into account lidar hits and pass-throughs for each 3D grid element (voxel). Standard occupancy mapping methods such as OHM assume lidar poses are perfectly known. However, in practice, poses are estimated by other state-estimation systems (e.g., SLAM, pose-graph optimisation, GNSS) which are inevitably subject to estimation errors due to sensor noise. Ignoring pose estimation error leads to inaccurate maps. We address this challenge by incorporating the estimation uncertainty of the aerial vehicle’s trajectory from the pose-estimation system into occupancy mapping.
The occupancy mapping module estimates the occupancy probability of each voxel given aerial lidar measurements. We denote the map by $\{p(m_{i}\mid\mathcal{Z})\}_{i}$ where $m_{i}$ is a Bernoulli random variable indicating the occupancy of the $i$th voxel, and $\mathcal{Z}$ denotes the set of all measurements from the aerial vehicle (e.g. lidar, IMU, and GNSS). Voxel occupancy probability is
@@
where $\mathcal{X}$ denotes the robot’s trajectory (represented by a finite set of poses). The first integrand $p(m_{i}\,|\,\mathcal{X},\mathcal{Z})$ in (1) is the voxel occupancy probability computed by standard occupancy mapping systems given trajectory $\mathcal{X}$. Note that voxel occupancy $m_{i}$ is conditionally independent from sensor measurements given the trajectory $\mathcal{X}$. This term can be further simplified to $p(m_{i}\,|\,\mathcal{X},\mathcal{Z}_{\text{lidar}})$ where $\mathcal{Z}_{\text{lidar}}\subset\mathcal{Z}$ denotes the set of lidar measurements. The second integrand $p(\mathcal{X}\,|\,\mathcal{Z})$ is the posterior distribution over the trajectory given all measurements. We use Monte Carlo (MC) integration to approximate the intractable integral in (1) by sampling $N$ independent trajectories $\{\mathcal{X}^{[k]}\}_{k=1}^{N}$ from the trajectory posterior, where $\mathcal{X}^{[k]}\sim p(\mathcal{X}\mid\mathcal{Z})$, and computing the following expression,
$p(m_{i}\mid\mathcal{Z})\approx\frac{1}{N}\sum_{k=1}^{N}p(m_{i}\mid\mathcal{X}^{[k]},\mathcal{Z}_{\text{lidar}}).$ (2)
Equation (2) suggests the following scheme for incorporating pose-estimation uncertainty into occupancy mapping: (i) build an occupancy map using a standard occupancy mapping method such as OHM for each trajectory sample $\mathcal{X}^{[k]}$; and (ii) for each voxel compute the average of the resulting $N$ predicted occupancy probabilities. We term this approach Uncertainty-Aware (UA) occupancy mapping.
Note that the details of trajectory sampling depend in part on the underlying pose-estimation technique. For example, one may use ancestral sampling to sample a trajectory when the pose-estimation is obtained by integrating inertial measurements over time. If poses are estimated independently using absolute measurements (e.g., GPS), we just need to independently sample poses from the corresponding posteriors to obtain a trajectory sample. More complex pose-estimation systems such as SLAM typically provide a Gaussian approximation of pose marginals from which one can sample poses independently to obtain an approximate trajectory sample. Alternatively, one can avoid the error caused by breaking the correlated trajectory posterior into pose marginals by transporting samples generated from the standard normal distribution using the inverse Cholesky factor of the sparse information matrix provided by SLAM systems.
@@ IV-C Obstruction score estimation
Figure 3: We reason on free, occupied and uncertain voxels by integrating lidar measurements in an occupancy grid map. Later the occupancy probabilities are fused in our scoring method to estimate the final obstruction map.
We use our 3D UA-occupancy map and the estimated ground height to assign a simple obstruction risk score to each ground-level 2D cell. This score is computed for each vertical column of $n$ 3D voxels starting from the ground cell as illustrated in Fig. 3. Formally, the score function $b:\mathbb{Z}\times\mathbb{Z}\to[0,1]$ assigns a value between zero and one to each 2D cell (i.e., the grid $XY$ coordinates of the corresponding column in the world frame). A larger value of $b(s)$ for a cell $s\triangleq(x,y)$ indicates a higher risk of obstruction at that cell. The number $n$ of column voxels is chosen according to the clearance height ($h$ in Fig. 3) required by the robot to pass through a column. Hence, taller robots require a larger $n$.
A cell $s$ might be considered as obstructed if at least one of the voxels in the corresponding column is occupied. This choice leads to $b(s)=1-\prod_{i\in\mathcal{I}_{s}}p(m_{i}=0\mid\mathcal{Z})$. However, because of the complexity of forest environments and the sparsity of aerial lidar rays reaching the ground level, the resulting score function will likely fail to correctly capture traversable cells where only the first voxel above ground is occupied by grass or an overhanging foliage is occupying the top voxel in the stack. An effective heuristic estimate is to instead use the weighted average of voxel occupancy probabilities within a column as the corresponding obstruction score, where the (fixed) weights reflect the significance of voxel occupancy. The resulting score function $b$ is given by
@@
where $\mathcal{I}_{s}$ denotes the set of indices of the $n$ voxels that are part of the vertical stack at cell $s$, and $p(m_{i}=1\mid\mathcal{Z})$ is the UA-occupancy probability of voxel $i$ given the aerial measurements. Specifically, in our work $n=4$ which allows for a robot height of 111\,1m using a 0.250.250.25\,0.25m voxel size, and the voxels weights $w_{i}$ from bottom to top are set to $(1,2,2,2)$. The rationale behind these values is to assign a smaller weight to the first voxel above ground level since the tracked mobile robot used in our experiments can better navigate obstacles immediately above the ground (e.g., small rocks, grass).